OpenStudy (anonymous):

Solve: negative 1 over 2 x + 3 = –x + 7

pooja195 (pooja195):

-1/2x+3=-x+7 like that?

OpenStudy (anonymous):

ya

OpenStudy (driftracer305):

\[-\left(\begin{matrix}1 \\ 2x\end{matrix}\right) + 3\] or \[-\left(\begin{matrix}1 \\ 2x+3\end{matrix}\right)\]

OpenStudy (driftracer305):

@helpplease101101??

OpenStudy (denisaboichuk):

second one